find 200 sum k=100^k



1
Expert's answer
2021-09-01T07:34:02-0400
"\displaystyle\sum_{k=100}^{200}k=\displaystyle\sum_{k=1}^{200}k-\displaystyle\sum_{k=1}^{99}k"

"=\dfrac{200(200+1)}{2}-\dfrac{99(99+1)}{2}="

"=50(402-99)=15150"